Nancy can mow a yard in 3 hours and Peter can mow it in 4 hours. Suppose they both work together, how long would it take to finish?
Answer by nerdybill(7384)

Nancy can mow a yard in 3 hours and Peter can mow it in 4 hours. Suppose they both work together, how long would it take to finish?
.
Let x = time (hours) it takes for both
then
x(1/3 + 1/4) = 1
multiplying both sides by 12:
x(4 + 3) = 12
x(7) = 12
x = 12/7 hours
or
x = 1 hour and 43 minutes